OpenStudy (anonymous):

The capacities of two trucks are 3 tons and 4 tons respectively. If the smaller truck makes 18 more trips than the larger, it can deliver 12 more tons of freight than the larger. How many trips does each truck make?

OpenStudy (saifoo.khan):

Let x represent the number of trips that the larger truck makes. : x = no. of trips of the larger truck then (x+18) = no. of trips the smaller truck : Small truck tonnage - large truck tonnage = 12 tons 3(x+18) - 4x = 12 3x + 54 - 4x = 12 3x - 4x = 12 - 54 -x = -42 x = 42 trips by the large truck then 42 + 18 = 60 trips by the small truck

OpenStudy (anonymous):

THANK YOU!!!! :D

OpenStudy (saifoo.khan):

uw